With only the small valve open, all of the liquid can be drained from a large vat in 4 hours. With only the large valve open, all of the liquid can be drained from the same vat in 2 hours. How long would it take to drain the vat with both valves open?
Answer by Nate(3500)

Rate of small valve: 1/4
Rate of large valve: 1/2
Rate * Time = Work
t/4 + t/2 = 1
4t/4 + 4t/2 = 4
t + 2t = 4
3t = 4
t = 4/3
1 hour and 20 minutes